Imagine that the Mini-Dow Average (MDA) is based on the closing prices of 5 stocks. The divisor used in the calculation of the MDA is currently 0.765. The closing prices for each of the 5 stocks in the MDA today and exactly 1 year ago, when the divisor was 0.780, are given in the following table: EEB a. Calculate the MDA today and that of a year ago b. Compare the values of the MDA calculated in part a and describe the apparent market behavior over the last year. Was it a bull or bear market? a. The MDA today is (Round to two decimal places.) The MDA one year ago was(Round to two decimal places.) b. Compare the values of the MDA calculated in part (a) and describe the apparent market behavior over the last year. Was it a bull or bear market? (Select the best answer below.) Data Table O A. It was a bear market. O B. It was a bull market. Stock Ace Computers Coburn Motor Company National Soap & Cosmetics Ronto Foods Wings Aircraft (Click on the icon located on the top-right corner of the data table below in order to copy its contents into a spreadsheet.) Today 73.53 37.96 109.89 71.63 97.31 1 Year Ago $64.99 34.34 96.95 72.03 88.87 Print Done
Answer #1

MDA today = sum of prices today/ Divisor

= (73.53+37.96+109.89+ 71.63 + 97.31) / 0.765=510.22

MDA one year before = Sum of prices/ Divisor

= (64.99+ 34.34+96.95+ 72.03+ 88.87)/0.78

=457.92

Since the Index has increased, it is a bull market